OpenStudy (anonymous):

sin^-1 [(radical 3)/2]

OpenStudy (anonymous):

\[\text{ArcSin}\left[\frac{\sqrt{3}}{2}\right]=\frac{\pi }{3}=60{}^{\circ} \]

OpenStudy (anonymous):

how did u get pi over 3?

OpenStudy (anonymous):